Can some help me on this question of probability? Please show your work.
For this question pretend you are drawing cards without replacement from the infamous "Iraq's Most Wanted" deck issued by the U.S. Military. If you are drawing from the full deck of 55 cards, what are the following probabilities:
a. You draw a card that is not Saddam Hussein
b. You draw three cards, which end up being Saddam Hussein and his two sons.
c. You draw 14 cards and not one of them is Saddam Hussein.

Answer by longjonsilver(2297)

55 cards? A deck of cards have 52 (+ 2 jokers).
Using a standard deck...
1. P(a certain card) = 1/52
2. P(3 certain cards, ABC) = P(1st is one of the 3) + P(2nd is one of the other 2) + P(3rd is the last one)
P(ABC) = 3/52 * 2/51 * 1/50 = 1/22100
3. P(not drawing card A in 14 attempts) = longwinded way, so you can see...
51/52 * 50/51 * 49/50 * 48/49 * ... * 38/39.
cancel out all the duplicate numbers to leave 38/52 = 0.731 to 3dp.

jon.